OpenStudy (anonymous):

you roll a number cube twice. Find each probability. 1. P(6 then 3) 2. P(odd then even)

OpenStudy (anonymous):

\[\frac{1}{6}\times \frac{1}{6}\] for the first one

OpenStudy (anonymous):

\[\frac{1}{2}\times \frac{1}{2}\] for the second one

OpenStudy (anonymous):

assuming of course that by "number cube" they mean dice and that the dice are fair. number cube?? indeed